OpenStudy (anonymous):

(In Proportion Form)...60% of 10 is what number and what number is 25% of 160... please help? Thanks! =)

OpenStudy (anonymous):

10/0.6=50/3 160*0.25=40

jhonyy9 (jhonyy9):

10 --- 100% x ---- 60% ------------- x=(60*10)/100 = 600/100 = 6 x=6 160 --- 100% x ----- 25% --------------- x=(25*160)/100 = 160/4 = 40 so 25% of 160 =40